The length of a rectangle has the double of its wide more four meters. If the total area is 336 m2 , What are its measures? please, would you like help me? .

Answer by stanbon(75887)

The length of a rectangle has the double of its wide more four meters. If the total area is 336 m2 , What are its measures?
-----------
Let the width be "x".
Then the length is 2w+4.
----
Area = length*width
336 = (2w+4)*w
2w^2 + 4w - 336 = 0
----
w^2 + 2w - 168 = 0
---
Factor:
(w+14)(w-12) = 0
---
Usable Answer:
width = 12
length = 2w+4 = 28
-----------------------------
Cheers,
Stan H.
=============